OpenStudy (anonymous):

help solve for x? please! : 13(x - 3) = 39

jimthompson5910 (jim_thompson5910):

13(x - 3) = 39 x - 3 = 39/13 x-3 = 3 Keep going and tell me what you get

OpenStudy (anonymous):

x=6

jimthompson5910 (jim_thompson5910):

you got it

jimthompson5910 (jim_thompson5910):

Check: 13(x - 3) = 39 13(6 - 3) = 39 13(3) = 39 39 = 39 answer is verified